OpenStudy (anonymous):

(3 × 3)(5 – 3)(6 + 2) = x2. What’s the value of x?

OpenStudy (anonymous):

\[ (3 × 3)(5 – 3)(6 + 2) = x^{2}\] \[\sqrt{(9)(2)(8)} =\sqrt{x^{2} }\] x=12